OpenStudy (anonymous):

In an isosceles triangle whose perimeter is 54 centimeters, the base is x centimeters and the measure of each leg is y centimeters. a. Write an equation that expresses the relationship between x and y. b. Find two ordered pairs in the solution set of the equation that you wrote in part A.

OpenStudy (anonymous):



OpenStudy (anonymous):

x+y+y= 54 x+2y= 54

OpenStudy (anonymous):

is that the answer for part A or part B

OpenStudy (anonymous):

no the above anwer is onlt for your question A

OpenStudy (anonymous):

so what the answer for question B?

OpenStudy (anonymous):

Question B, i have no idea. sorry ..

OpenStudy (anonymous):

oh okay..

OpenStudy (anonymous):

by the way. For part A , am i supposed to divide?

OpenStudy (anonymous):

i am trying for anwer, if i got i will let u know.. okk?

OpenStudy (anonymous):

yes sure.

OpenStudy (anonymous):

y u think to ivide??

OpenStudy (anonymous):

because you can divide 54 by 2

OpenStudy (anonymous):

here we cant divide. Becase that is not 2, it is " 2y.". it is a variable

OpenStudy (anonymous):

ooohh okay :)

OpenStudy (anonymous):

oky

OpenStudy (anonymous):

(10,22) solve the equation

OpenStudy (anonymous):

(10,22 ) is answer for part B

OpenStudy (anonymous):

let substitute the above values in place of x and y , then you will get 10-2*22= 54 ie(10, 22) pair solve the equation

OpenStudy (anonymous):

Your question is for 2 ordered pairs, so answers are (10,22) and ( 6, 24)

OpenStudy (anonymous):

hope you understand..